# Critical Quantum Sensing: a tutorial on parameter estimation near quantum phase transitions

### v1

@@ VI.7 Accessible measurements
Thus far, our discussion has revolved around the optimal attainable precision—as quantified by the QFI. This assumes that we are performing measurements in the optimal basis, allowing us to extract maximum information about the parameter of interest $\omega$ and lower bound the precision through the QCRB. However, the optimal measurement might not always be a practically implementable observable, and it is also important to consider more feasible measurement strategies. The LZ model is a simple two-level system, and therefore the choice of measurement basis is highly restricted: consisting only of two orthonormal Bloch vectors.
A natural choice is the basis formed by the spin-up $|\!\uparrow\rangle$ and spin-down $|\!\downarrow\rangle$ states, corresponding to the eigenstates of the $\hat{\sigma}_{z}$ operator. We will consider the metrological utility of this choice of observable and analyse its ensuing metrological capabilities at both zero and finite temperature to draw a comparison with the previous Sec. VI.2 and Sec. VI.6 where we considered the optimal sensitivity. We first consider the zero-temperature scenario. For this choice of observable, we obtain on average a measurement signal of
@@
$S_{\omega}=\frac{\left(\partial_{\omega}\langle\hat{\sigma}_{z}\rangle\right)^{2}}{\big(\Delta\langle\hat{\sigma}_{z}\rangle\big)^{2}}=\frac{g^{2}}{\left(g^{2}+\omega^{2}\right)^{2}}\;.$ (92)
Interestingly, the SNR provides an equivalent precision to the optimal QFI, $S_{\omega}=\mathcal{I}_{\omega}$. This implies that the transverse magnetization constitutes an optimal measurement at zero-temperature. Physically, we can understand that at $T=0$, the Bloch vector points around $\vec{n}=(g,0,\omega)/\Delta$ so changing $\omega$ tilts the state within the x{x}-z{z} plane, and $\hat{\sigma}_{z}$ is capable of reading that out with maximal efficiency.
A natural question to ask is if such a measurement $\langle\hat{\sigma}_{z}\rangle$ remains optimal when thermal fluctuations are introduced. As before, our probe state is now described by the canonical Gibbs state $\hat{\varrho}=e^{\hat{H}/T}/\mathcal{Z}$ as opposed to directly measuring the groundstate. Measuring the transverse magnetization at finite temperature we obtain on average a measurement signal of
@@
with $\Delta=\sqrt{\omega^{2}+g^{2}}$ the energy gap. Although not immediately obvious from the above expression, the resulting SNR is lower-bounded by the QFI of the thermal state
@@
This means that through the inclusion of thermal fluctuations, the transverse magnetization no longer constitutes an optimal measurement—despite being optimal at zero-temperature. This highlights the subtle nature of how external factors may augment precision measurement and how strategies which yield optimal precision at zero-temperature for example do not necessarily extend to finite temperature scenarios. In this instance, the optimal measurement is one which can fully exploit the population imbalance and mixed-state coherences.
The transverse magnetization suffers in this regard, as it is not capable of fully accessing the geometric component of mixed states when thermal fluctuations, thereby becoming a sub-optimal measurement in this regime. This can be seen explicitly in Fig. 9 where we compare the SNR to the optimal QFI. At $T\!\to\!0$ when the system is described by the groundstate and at $T\!\to\!\infty$ when the thermal state becomes maximally mixed the precision of both strategies is comparable. In contrast, at intermediate regime where the QFI becomes enhanced, the transverse magnetization is not capable of fully exploiting the features necessary to attain optimal precision.
Figure 9: LZ model comparison of optimal sensitivity to experimentally accessible measurement at finite temperature. The QFI $\mathcal{I}_{\omega}^{\beta}$ is shown by the dotted yellow line. The SNR for measuring in the $\hat{\sigma}_{z}$ basis is denoted by the black solid line. The latter constitutes a sub-optimal measurement, not capable of fully exploiting the thermal fluctuations. The frequency is set $\omega=1$ in both.
On a more general footing, there is an important caveat we must consider when choosing $\hat{\sigma}_{z}$ as our choice observable. We are assuming we have the freedom to choose the measurement basis. However, in many practical scenarios such freedom is not readily available. Often-times the only viable option is to measure in the energy eigenbasis. At first glance this seems problematic as in the energy eigenbasis the Hamiltonian becomes
@@
This implies that the signal of the population difference measurement is always constant—regardless of the parameters. Equivalently, this is analogous to stating the system is simply in its groundstate which offers no direct information about the parameters. In such instances, one can instead perform Ramsey interferometry which measures the energy separation between the two levels. In the rotating frame of the driving laser and following the Ramsey sequence, the signal becomes
@@
The SNR is maximized when $g=0$, which stands in contrast to the optimal condition derived from Eq. (75). This discrepancy highlights again an important point: for few-body systems, critical quantum metrology does not inherently provide any enhancement over standard quantum limit. These results will serve as a baseline for the following sections, where we explore many-body quantum systems that exhibit genuine quantum phase transitions.
@@ VI.9 Effective parameter in multiparameter estimation
with the rescaled Hamiltonian $\hat{H}_{\text{LZ}}^{\prime}$ being dimensionless and $\omega$ being treated as a unit of energy. Correspondingly, we can deduce that the system can be fully specified by the ratio of parameters $\Omega=g/\omega$. It then becomes obvious that $g$ and $\omega$ cannot be independently estimated from pure states. This is because no measurements using such a state can distinguish between different combinations of the parameters, $(\omega,g)$ and $(2\omega,2g)$ for example, as the ratio of the parameters is the same in both instances, and the state depends only on the ratio. Therefore, while true multiparameter estimation is prohibited, the appropriate object which can be estimated in this instance is the single effective parameter $\Omega=g/\omega$.
Using the rescaled Hamiltonian $\hat{H}_{\text{LZ}}^{\prime}$ the effective single parameter QFI can be computed for the groundstate yielding
@@
where in the second equality we express the effective QFI in terms of the original bare parameters. In the more general case, discerning the nature of effective parameters and their corresponding precision is not as straightforward. This is especially true in complicated many-body systems and in systems undergoing phase transitions. In the latter case, parameter dependencies may naturally emerge in the vicinity of critical points, where universal physics is encoded in order parameters that combine the bare Hamiltonian parameters in non-trivial ways. Moreover, emergent Hamiltonian symmetries at criticality—not present away from the transition, may impose additional constraints on the system, potentially giving rise to such metrological symmetries.
@@
Given that it may not always be possible to discern the effective parameter encoding and hence attainable precision, from the quantum state alone in complicated many-body critical systems, a more systemic approach is desirable.
@@ VI.10 Deducing effective parameters in multiparameter estimation
which provides a non-zero constraint $\hat{\mathcal{I}}(\vec{\theta})\vec{u}_{b}=G_{b}\vec{u}_{b}$ that can be utilized to discern the effective parameter encoding $\Omega=f(\vec{\theta})$ by inverting the set of non-linear partial differential equations from the condition $\vec{u}_{b}=\vec{\nabla}_{\theta}\Omega/|\vec{\nabla}_{\theta}\Omega|$ in the case of a single zero eigenvalue. By solving the set of partial differential equations we obtain the following solutions for the effective QFI for the LZ model
@@
where only the first solution is physical. The above demonstrates that care is essential when deducing the effective parameter encoding, as even in simple cases, unphysical solutions may arise. The presence of multiple solutions also implies potential branch-cut issues when performing numerical computations. Once the functional form of the effective parameter, $\Omega=g/\omega$, has been deduced from the singular QFIM one can, in principle, one can reparametrize the Hamiltonian or groundstate accordingly to find the corresponding effective QFI as before in Eq. (103). For situations where such a reparametrization is not possible, an alternative is required.
@@ VI.11 Bounding precision of effective parameters in multiparameter estimation
A widely adopted approach to deal with singular multiparameter estimation and to formulate approximate bounds on the precision is to apply the Moore-Penrose pseudoinverse to the QFIM [119, 168, 123, 171, 167]. The pseudoinverse of the QFIM $\left[\hat{\mathcal{I}}(\vec{\theta})\right]^{+}$ always exists and is well defined. The interpretation of the pseuodinverse as a lower bound $\text{Cov}[\vec{\theta}]\geq\left[\mathcal{M}\hat{\mathcal{I}}(\vec{\theta})\right]^{+}$ on the covariance matrix of parameters is known to be overly optimistic compared to the true QCRB [171]. The pseudoinverse of the LZ model can readily be computed in this case
$\left[\hat{\mathcal{I}}_{\text{LZ}}\right]^{+}=\begin{pmatrix}g^{2}&\omega g\\ \omega g&\omega^{2}\end{pmatrix}\;$. (109)
We can then utilize the previously deduced effective parameter encoding $\Omega=g/\omega$ in conjunction with the pseudoinverse Eq. (109) to deduce the true precision of the effective parameter, as opposed to the optimistic one arising from $\text{Cov}[\vec{\theta}]\geq\left[\mathcal{M}\hat{\mathcal{I}}(\vec{\theta})\right]^{+}$ [121]. This can be achieved by using the linearized error propagation formula
@@
with $\vec{\Omega}$ the vector of effective parameters and $\hat{J}=\nabla\Omega(\omega,g)=(\partial_{\omega}\Omega,\partial_{g}\Omega)\equiv(1/g,-\omega/g^{2})$ the Jacobian. We obtain the following precision from substitution of the above,
@@
where we treat the values $\text{Var}(\omega)$, $\text{Var}(g)$, and $\text{Cov}(\omega,g)$ as elements of the pseudoinverse $\left[\mathcal{M}\hat{\mathcal{I}}_{\text{LZ}}\right]^{+}$. From this the precision of $\Omega$ can be found to
$\text{Var}(\Omega)=\frac{(g^{2}-\omega^{2})^{2}}{\omega^{4}}\;,$ (112)
whose inverse is precisely the effective QFI found by using the reparametrized Hamiltonian $\hat{H}_{\text{LZ}}^{\prime}$ in Eq. (103). We can therefore understand that discerning the effective parameter encoding is crucial in the pseudoinverse approach to retrieve the exact saturable bound, as opposed to the optimistic one.
@@ VI.12 Finite temperature multiparameter estimation
In the context of single parameter estimation we have seen in Sec. VI.6 how finite temperature fluctuations may serve to further enhance the attainable precision beyond what is achievable with the groundstate alone [149, 150, 55]. This improvement arises from two effects: additional information encoded in thermal populations and the mixing of states in the statistical ensemble $\hat{\varrho}=\sum_{n}e_{n}|e_{n}\rangle\langle e_{n}|$. However, this benefit is limited to low but finite temperatures: in the limit $T\to 0$ the precision is governed by the groundstate, while at $T\to\infty$ the state becomes maximally mixed and no useful information can be recovered.
@@
Figure 11: Multiparameter estimation at finite temperature for the LZ model. (a) Total finite-temperature QFI for estimating $\omega$ in the single parameter case when the control parameter $g$ is assumed to be known. (b) Total finite-temperature QFI for the joint estimation of $\omega$ and $g$ in the multiparameter case. Here the control parameter $g$ is also treated as an unknown parameter. (c) Comparison of the single and multiparameter QFI shown in the black and dashed orange lines respectively. The frequency is fixed $\omega=1$ in all panels; $g=0.8$ for panel (c).
To lift the QFIM singularity and allow for the estimation of our target $\omega$ as an independent parameter, one must break the metrological symmetry. This serves to remove the constraints rendering the encoding redundant and restores full rank to the QFIM. One method of achieving this is to add known external controls on the system [66]. Another possibility is to leverage thermal fluctuations: by treating temperature as a known quantity, it may in some circumstances serve as a mechanism for breaking the metrological symmetry, decoupling the parameter dependencies. We illustrate this point by considering the estimation of splitting $\omega$ and coupling $g$ in the LZ model, but this time at finite temperature. The QFIM is written in the following basis
@@
Introducing temperature as a known parameter, the previous constraint rendering the QFIM singular for the groundstate is broken and the QFIM becomes invertible with $\text{det}\left[\hat{\mathcal{I}}_{\text{LZ}}^{\beta}\right]=\mathcal{I}_{\omega\omega}^{\beta}\mathcal{I}_{gg}^{\beta}-\big(\mathcal{I}_{g\omega}^{\beta}\big)^{2}\neq 0$. The determinant is given by
@@
The above expression reveals an interesting competition: the denominator $T^{2}\Delta^{2}$ suggests a divergence in the determinant as the energy gap closes $\Delta\!\to\!0$ or temperature vanishes $T\!\to\!0$, potentially indicating sensitivity is restored in this regime compared to the singular groundstate solution. However, this is counteracted by the numerator, specifically $\text{csch}^{2}\big(\Delta/T\big)$ term, which strongly suppresses sensitivity at low temperatures ($T\!\to\!0$), recovering the groundstate result of a singular QFIM. The competition between these two terms determines the overall behaviour of the determinant. In particular, when $T\ll\Delta$ and $|g|\gg\omega$ the system is described by its groundstate properties and the determinant vanishes. In this region the thermal population of the excited states are exponentially suppressed. Conversely in the opposite limit, $T\to\infty$, the thermal state becomes maximally mixed and no information about the parameters is encoded, and the QFIM becomes singular.
Between these extremes lies an optimal regime, $T\sim\Delta$ and $|g|\sim\omega$, for which thermal fluctuations break the metrological symmetry and restores full-rank to the QFIM. Therefore, in the sense of multiparameter estimation, metrological utility is enhanced in this region by making estimation of our target $\omega$ permissible when the control parameter $g$ is unknown. This result highlights a counter-intuitive yet powerful principle: under certain conditions, thermal noise can become a valuable resource in quantum metrology by unlocking otherwise inaccessible parameter information.
VI.13 Finite temperature precision comparison of single and multiparameter cases
Having shown that thermal fluctuations can break the metrological symmetry and restore full rank to the QFIM within a well-defined regime, we are now in a position to tolerate uncertainty in the control parameter $g$. This allows us to meaningfully compare single and multiparameter estimation strategies for our target parameter $\omega$. In the multiparameter setting, the relevant figure of merit for estimation precision is the inverse QFIM,

### v2

@@ VI.7 Accessible measurements
Thus far, our discussion has revolved around the optimal attainable precision—as quantified by the QFI. This assumes that we are performing measurements in the optimal basis, allowing us to extract maximum information about the parameter of interest $\omega$ and lower bound the precision through the QCRB. However, the optimal measurement might not always be a practically implementable observable, making it important to also consider more feasible measurement strategies. The LZ model is a simple two-level system, and therefore the choice of measurement basis is highly restricted: it consists only of two orthonormal Bloch vectors.
A natural choice of measurement basis is that formed by the spin-up $|\!\uparrow\rangle$ and spin-down $|\!\downarrow\rangle$ states, which corresponds to the eigenstates of the $\hat{\sigma}_{z}$ operator. We now examine the metrological performance associated with this observable, analysing its sensitivity at both zero and finite temperature. This allows for a direct comparison with Sections VI.2 and VI.6, where we considered the optimal sensitivity. We first examine the zero-temperature scenario. For this choice of observable, we obtain on average a measurement signal of
@@
$S_{\omega}=\frac{\left(\partial_{\omega}\langle\hat{\sigma}_{z}\rangle\right)^{2}}{\left(\Delta\hat{\sigma}_{z}\right)^{2}}=\frac{g^{2}}{\left(g^{2}+\omega^{2}\right)^{2}}\;.$ (92)
Interestingly, the SNR provides the same precision as the QFI, $S_{\omega}=\mathcal{I}_{\omega}$. This implies that the transverse magnetization constitutes an optimal measurement at zero-temperature. Physically, we can understand that at $T=0$, the Bloch vector points around $\vec{n}=(g,0,\omega)/\Delta$ so changing $\omega$ tilts the state within the x{x}-z{z} plane, and $\hat{\sigma}_{z}$ is capable of reading that out with maximal efficiency.
A natural question to ask is if such a measurement remains optimal when thermal fluctuations are introduced. As before, the probe state is now described by the canonical Gibbs state $\hat{\varrho}=e^{-\hat{H}/T}/\mathcal{Z}$, rather than a pure ground state as in the zero-temperature case. By measuring the transverse magnetization at finite temperature, we obtain on average a measurement signal of
@@
where $\Delta=\sqrt{\omega^{2}+g^{2}}$ is the energy gap. Although not immediately obvious from the above expression, the resulting SNR is lower-bounded by the QFI of the thermal state
@@
This means that through the inclusion of thermal fluctuations, the transverse magnetization no longer constitutes an optimal measurement—despite being optimal at zero-temperature. This highlights the subtle nature of how external factors can influence precision measurements, and that strategies yielding optimal performance at zero temperature do not necessarily extend to finite-temperature regimes. In this case, the optimal measurement is one that fully exploits both the population imbalance and the coherences present in the mixed state.
The transverse magnetization suffers in this regard, as it cannot fully access the geometric component of mixed states in the presence of thermal fluctuations, thereby becoming a suboptimal measurement in this regime. This can be seen explicitly in Fig. 9, where we compare the SNR to the QFI. For $T\!\to\!0$, when the system is described by the ground state, and for $T\!\to\!\infty$, when the thermal state becomes maximally mixed, the precision of both strategies is comparable. In contrast, in the intermediate regime, where the QFI is enhanced, the transverse magnetization is not capable of fully exploiting the features necessary to attain optimal precision.
Figure 9: Comparison of optimal and experimentally accessible sensitivities in the LZ model at finite temperature. The QFI $\mathcal{I}_{\omega}^{\beta}$ is shown by the dotted yellow line. The SNR obtained from measurements in the $\hat{\sigma}_{z}$ basis is shown by the black solid line. The latter constitutes a suboptimal measurement, not capable of fully exploiting the thermal fluctuations. In both cases, the frequency is set to $\omega=1$.
On a more general footing, an important caveat arises when adopting $\hat{\sigma}_{z}$ as the measurement observable. We are assuming we have the freedom to choose the measurement basis. However, in many practical scenarios such freedom is not readily available. Often, the only viable option is to measure in the energy eigenbasis. At first glance, this seems problematic since, in that basis, the Hamiltonian takes the diagonal form
@@
This implies that the population-difference signal is always constant—regardless of the parameters values. Equivalently, this is analogous to stating the system is simply in its ground state, which offers no direct information about the parameters. In such situations, one can instead perform Ramsey interferometry, which measures the energy separation between the two levels. In the rotating frame of the driving laser and following the Ramsey sequence, the signal becomes
@@
The SNR is maximized when $g=0$, which stands in contrast to the optimal condition derived from Eq. (75). This discrepancy highlights again an important point: for few-body systems, critical quantum metrology does not inherently provide any enhancement over the standard quantum limit. These results will serve as a baseline for the following sections, where we explore many-body quantum systems that exhibit genuine quantum phase transitions.
@@ VI.9 Effective parameter in multiparameter estimation
with the rescaled Hamiltonian $\hat{H}_{\text{LZ}}^{\prime}$ being dimensionless and $\omega$ being treated as a unit of energy. Correspondingly, we can deduce that the system can be fully specified by the ratio of parameters $\Omega=g/\omega$. It then becomes obvious that $g$ and $\omega$ cannot be independently estimated from pure states. This is because no measurements using such a state can distinguish between different combinations of the parameters, $(\omega,g)$ and $(2\omega,2g)$ for example, as the ratio of the parameters is the same in both instances, and the state depends only on this ratio. Therefore, while true multiparameter estimation is prohibited, the appropriate object which can be estimated in this case is the single effective parameter $\Omega=g/\omega$.
Using the rescaled Hamiltonian $\hat{H}_{\text{LZ}}^{\prime}$, the effective single parameter QFI can be computed for the ground state, yielding
@@
where in the second equality we express the effective QFI in terms of the original bare parameters. In the more general case, discerning the nature of effective parameters and their corresponding precision is not as straightforward. This is especially true in complicated many-body systems and in systems undergoing phase transitions. In the latter case, parameter dependencies may naturally emerge in the vicinity of critical points, where universal physics is encoded in order parameters that combine the bare Hamiltonian parameters in non-trivial ways. Moreover, emergent Hamiltonian symmetries at criticality—which are absent away from the transition—may impose additional constraints on the system, potentially giving rise to such metrological symmetries.
@@
Since it is not always possible to discern the effective parameter encoding, and thus the attainable precision, from the quantum state alone in complicated many-body critical systems, a more systematic approach is desirable.
@@ VI.10 Deducing effective parameters in multiparameter estimation
which provides a non-zero constraint $\hat{\mathcal{I}}(\vec{\theta})\vec{u}_{b}=G_{b}\vec{u}_{b}$ that can be utilized to discern the effective parameter encoding $\Omega=f(\vec{\theta})$ by inverting the set of non-linear partial differential equations, derived from the condition $\vec{u}_{b}=\vec{\nabla}_{\theta}\Omega/|\vec{\nabla}_{\theta}\Omega|$ in the case of a single zero eigenvalue. By solving the set of partial differential equations, we obtain the following solutions for the effective QFI for the LZ model:
@@
where only the first solution is physical. The above demonstrates that care is essential when deducing the effective parameter encoding, as even in simple cases, unphysical solutions may arise. The presence of multiple solutions also implies potential branch-cut issues when performing numerical computations. Once the functional form of the effective parameter, $\Omega=g/\omega$, has been deduced from the singular QFIM one can, in principle, reparametrize the Hamiltonian or ground state accordingly to find the corresponding effective QFI, as done before in Eq. (103). For situations where such a reparametrization is not possible, an alternative is required.
@@ VI.11 Bounding precision of effective parameters in multiparameter estimation
A widely adopted approach to deal with singular multiparameter estimation and to formulate approximate bounds on the precision is to apply the Moore-Penrose pseudoinverse to the QFIM [131, 181, 134, 184, 180]. The pseudoinverse of the QFIM $\left[\hat{\mathcal{I}}(\vec{\theta})\right]^{+}$ always exists and is well defined. The interpretation of the pseuodinverse as a lower bound $\text{Cov}[\vec{\theta}]\geq\left[\hat{\mathcal{I}}(\vec{\theta})\right]^{+}$ on the covariance matrix of parameters is known to be overly optimistic compared to the true QCRB [184]. The pseudoinverse of the LZ model can readily be computed in this case
$\left[\hat{\mathcal{I}}_{\text{LZ}}\right]^{+}=\begin{pmatrix}g^{2}&-\omega g\\ -\omega g&\omega^{2}\end{pmatrix}\;$. (109)
We can then utilize the previously deduced effective parameter encoding $\Omega=g/\omega$ in conjunction with the pseudoinverse in Eq. (109) to deduce the true precision of the effective parameter, as opposed to the optimistic one arising from $\text{Cov}[\vec{\theta}]\geq\left[\mathcal{M}\hat{\mathcal{I}}(\vec{\theta})\right]^{+}$ [33]. This can be achieved by using the linearized error propagation formula
@@
with $\vec{\Omega}$ the vector of effective parameters and $\hat{J}=\nabla\Omega(g,\omega)=(\partial_{g}\Omega,\partial_{\omega}\Omega)\equiv(1/\omega,-g/\omega^{2})$ the Jacobian. We obtain the following precision from substitution of the above
@@
where we treat the values $\text{Var}(\omega)$, $\text{Var}(g)$, and $\text{Cov}(\omega,g)$ as elements of the pseudoinverse $\left[\hat{\mathcal{I}}_{\text{LZ}}\right]^{+}$. From this, the precision of $\Omega$ can be found to be
$\text{Var}(\Omega)=\frac{(g^{2}+\omega^{2})^{2}}{\omega^{4}}\;,$ (112)
whose inverse is precisely the effective QFI found by using the reparametrized Hamiltonian $\hat{H}_{\text{LZ}}^{\prime}$ in Eq. (103). We can therefore understand that discerning the effective parameter encoding is crucial in the pseudoinverse approach to retrieve the exact saturable bound, as opposed to the optimistic one [33].
@@ VI.12 Finite temperature multiparameter estimation
In the context of single parameter estimation, we have seen in Section VI.6 how finite temperature fluctuations may serve to further enhance the attainable precision beyond what is achievable with the ground state alone [162, 163, 59]. This improvement arises from two effects: additional information encoded in thermal populations and the mixing of states in the statistical ensemble $\hat{\varrho}=\sum_{n}e_{n}|e_{n}\rangle\langle e_{n}|$. However, this benefit is limited to low but finite temperatures: in the limit $T\to 0$ the precision is governed by the ground state, while at $T\to\infty$ the state becomes maximally mixed and no useful information can be recovered.
@@
Figure 11: Multiparameter estimation at finite temperature for the LZ model. (a) Total finite-temperature QFI (89) for estimating $\omega$ in the single parameter case when the control parameter $g$ is assumed to be known. (b) Total finite-temperature QFI for the joint estimation of $\omega$ and $g$ in the multiparameter case. Here the control parameter $g$ is also treated as an unknown parameter. (c) Comparison of the single and multiparameter QFI shown in the black and dashed orange lines respectively. In all panels we set $\omega=1$; $g=0.8$ for panel (c).
To lift the QFIM singularity and allow for the estimation of our target $\omega$ as an independent parameter, one must break the metrological symmetry. This serves to eliminate the constraints that render the encoding redundant, thereby restoring full rank to the QFIM. One method of achieving this is to add known external controls on the system [32]. Another possibility is to leverage thermal fluctuations: by treating temperature as a known quantity, it may in some circumstances serve as a mechanism for breaking the metrological symmetry, decoupling the parameter dependencies. We illustrate this point by considering the estimation of $\omega$ and $g$ in the LZ model, but this time at finite temperature $T$. In this case, the QFIM reads
@@
Introducing temperature as a known parameter, the previous constraint rendering the QFIM of the ground state singular is broken, and the QFIM becomes invertible with $\text{det}\left[\hat{\mathcal{I}}_{\text{LZ}}^{\beta}\right]=\mathcal{I}_{\omega\omega}^{\beta}\mathcal{I}_{gg}^{\beta}-\big(\mathcal{I}_{g\omega}^{\beta}\big)^{2}\neq 0$. Indeed, we have now
@@
The above expression reveals an interesting competition: the denominator $T^{2}\Delta^{2}$ suggests a divergence in the determinant as the energy gap closes, $\Delta\!\to\!0$, or temperature vanishes, $T\!\to\!0$, potentially indicating sensitivity is restored in this regime compared to the singular ground state solution. However, this is counteracted by the numerator, specifically by the $\text{csch}^{2}\big(\Delta/T\big)$ term, which strongly suppresses sensitivity at low temperatures ($T\!\to\!0$), recovering the ground state result of a singular QFIM. The competition between these two terms determines the overall behaviour of the determinant. In particular, when $T\ll\Delta$ and $|g|\gg\omega$ the system is described by its ground state properties and the determinant vanishes. In this region the thermal population of the excited states is exponentially suppressed. Conversely in the opposite limit, $T\to\infty$, the thermal state becomes maximally mixed and no information about the parameters is encoded, thus the QFIM becomes singular.
Between these extremes lies an optimal regime, $T\sim\Delta$ and $|g|\sim\omega$, for which thermal fluctuations break the metrological symmetry and restore full rank to the QFIM. Therefore, for multiparameter estimation, metrological utility is enhanced in this region by making the estimation of our target $\omega$ possible when the control parameter $g$ is unknown. This result highlights a counter-intuitive yet powerful principle: under certain conditions, thermal noise can become a valuable resource in quantum metrology by unlocking otherwise inaccessible parameter information.
VI.13 Comparison of precision at finite temperature for single and multiparameter cases
Having shown that thermal fluctuations can break the metrological symmetry and restore full rank to the QFIM within a well-defined regime, we are now in a position to tolerate uncertainty in the control parameter $g$. This allows us to meaningfully compare single and multiparameter estimation strategies for our target parameter $\omega$. In the multiparameter setting, the relevant figure of merit for the estimation precision is the inverse of the QFIM